OpenStudy (anonymous):

4(3-x) (4/3)power -5=59

OpenStudy (anonymous):

\[\LARGE 4(3-x) \cdot \left(\frac43\right)^{-5}=59\] ? :)

OpenStudy (anonymous):

aughh.. \[\LARGE 4(3-x)^{4/3}-5=59\] ?

OpenStudy (anonymous):

No, the 4/3 is after the (3-x). it's not being multiplied